# HybGRAG: Hybrid Retrieval-Augmented Generation on Textual and Relational Knowledge Bases

### v1

@@ 1 Introduction
Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Guu et al., 2020) enables large language models (LLMs) to access the information from an unstructured document database. This allows LLMs to address unknown facts and solve Open-Domain Question Answering (ODQA) with additional textual information. Graph RAG (GRAG) has extended this concept by retrieving information from structured knowledge bases, where documents are interconnected by relationships. Existing GRAG methods focus on two directions: extracting relational information from knowledge graphs (KGs) and leverages LLMs for Knowledge Base Question Answering (KBQA) (Yasunaga et al., 2021; Sun et al., 2024; Jin et al., 2024; Mavromatis and Karypis, 2024), and building relationships between documents in the database to improve ODQA performance (Li et al., 2024a; Dong et al., 2024; Edge et al., 2024).
@@ 3. Interpretable ?
In Table 1, HybGRAG is the only work that satisfies all the properties and solves HQA. In Fig. 2, evaluating in a HQA benchmark STaRK, HybGRAG outperforms the second-best baseline with relative improvements in Hit@$1$ 47% in STaRK-MAG, and 55% in STaRK-Prime, respectively.
@@ 2 Proposed Insights: Challenges in HQA
What are the new challenges in HQA over SKB that cannot be solved by the existing RAG and GRAG methods? In this section, we first define the problem, and then conduct experiments to uncover two critical insights in HQA, laying the foundation for designing our method for HQA.
@@ 2.1 Problem Definition
A semi-structured knowledge base (SKB) consists of a KG $G=(\mathcal{E},\mathcal{R})$, where $\mathcal{E}$ and $\mathcal{R}$ are sets of entities and relations, respectively, and a set of text documents $\mathcal{D}$. Entity and relation types are denoted as $\mathcal{T}_{E}$ and $\mathcal{T}_{R}$, respectively. Each hybrid question $q$ in SKB involves semi-structured information, namely, textual and relational information. We define hybrid question answering (HQA) as follows:
@@ 2.2 C1: Hybrid-Sourcing Question
To investigate whether it is necessary to leverage both textual and relational information to answer hybrid questions, we conduct an experiment to show that text documents and KG contain useful but non-overlapping information. As a retriever that uses only textual information, vector similarity search (VSS) (Karpukhin et al., 2020) performs retrieval and ranking by comparing the question and documents in the embedding space (“ada-002”); as a retriever that uses only relational information, Personalized PageRank (PPR) (Andersen et al., 2006) performs random walks starting from the topic entities identified by an LLM (Claude 3 Sonnet) and ranks neighboring entities based on their connectivity in KG of SKB.
@@ 3.1 Retriever Bank (for C1)
The hybrid retrieval module takes the identified topic entities $\hat{\mathcal{E}}$ and useful relations $\hat{\mathcal{R}}$ as input. It uses a graph retriever to extract entities in the ego-graph of $\hat{\mathcal{E}}$, connected by $\hat{\mathcal{R}}$. For example, in Fig. 1, {^=$\{\hat{\mathcal{E}}=\{\text{John Smith}\},\hat{\mathcal{R}}=\{\text{author writes paper}\}\}$ and the graph retriever extracts the entities/papers connected by the path “John Smith -> author writes paper -> {papers}”. If more than one ego-graph is extracted, their intersection is used as the final result. Finally, to solve hybrid questions, we propose ranking the documents associated with the extracted entities using a VSS ranker between question $q$ and documents $\mathcal{D}$. This ensures the synergy between the relational and textual information.
@@ 3.2 Critic Module (for C2)
To solve this, we propose the critic module, which provides feedback to help the router perform better question routing. Instead of using a single LLM to complete this complicated task, we divide the critic into two parts, an LLM validator $C_{val}$ to validate the correctness of the retrieval $\mathcal{X}$, and an LLM commentor $C_{com}$ to provide feedback $f_{t}$ if the retrieval is incorrect. This divide-and-conquer step, similar to previous works (Gao et al., 2022; Asai et al., 2024), is crucial to our critic module, offering two key advantages: (1) By breaking a difficult task into two easier ones, we can now leverage pre-trained LLMs to solve them while maintaining good performance. This resolves the issue when the labels are not available for fine-tuning an LLM critic. (2) Since the tasks of $C_{val}$ and $C_{com}$ are independent, they can each have their own exclusive contexts, preventing the inclusion of irrelevant information and avoiding the “lost in the middle” phenomenon (Shi et al., 2023; Liu et al., 2024).
@@
Commentor
The LLM commentor $C_{com}$ aims to provide feedback $f$ to help the router refine question routing. To effectively guide the router, we construct corrective feedback that it can easily understand. In more detail, it points out the error(s) in each action, such as incorrect identification of topic entities, as shown in Table 4. Unlike natural language feedback, which may cause uncertainty or inconsistency depending on the LLM used, our corrective feedback provides clear guidance on how to refine the question routing. Furthermore, it leverages in-context learning (ICL) to provide sophisticated feedback. We collect a small number of successful experiences ($\approx 30$) in the training set as examples, with each experience $\{s_{t},\hat{\mathcal{E}}_{t},\hat{\mathcal{R}}_{t},f_{t+1}\}$ comprising a pair of router action and feedback, which is verified by ground truth. During inference, the commentor gives high-quality feedback based on multiple pre-collected examples.
@@ 3.3 Overall Algorithm
The algorithm of HybGRAG is in Algo. 1. Given a question $q$, in iteration $t$, the router determines $s_{t}$, ^$\hat{\mathcal{E}}_{t}$ and ^$\hat{\mathcal{R}}_{t}$ to retrieve the references $\mathcal{X}_{t}$ from both $G$ and $\mathcal{D}$ in SKB, or only $\mathcal{D}$, with the selected retrieval module. The validator $C_{val}$ in the critic module then decides whether to accept $\mathcal{X}_{t}$ as the final answer or reject it. If $\mathcal{X}_{t}$ is rejected, the commentor $C_{com}$ generates feedback $f_{t+1}$ for the router to assist in refining its action in iteration $t+1$.
@@ 4 Experiments
    Effectiveness: How well does HybGRAG perform in real-world GRAG benchmarks?
@@
Figure 3: Design choices in HybGRAG are necessary in STaRK. We compare HybGRAG with two variants: a validator without validation context, and a commentor with only 5-shot. Oracle uses ground truth during inference.
@@
GRAG Benchmarks
We conduct experiments on two GRAG benchmarks: STaRK  [^1] Due to legal issue, only STaRK-MAG and STaRK-Prime are used in this article. (Wu et al., 2024b), which serves as the primary evaluation benchmark and focuses on retrieval, and CRAG (Yang et al., 2024), a complementary benchmark to evaluate end-to-end RAG performance. Detailed benchmark descriptions are provided in Appx. A.
@@ 4.1.2 Ablation Study (RQ2)
Critic Module
We compare HybGRAG variants with a validator without validation context, a commentor with only five shots, and those with oracles. The oracle has access to the ground truth, which gives the optimal judgement on the correctness of the output and the error type of question routing, if there is any. In Fig. 3, we show that HybGRAG performs the best with all our design choices, approaching the performance of an oracle.
Self-Reflection
In Figure 4, we demonstrate that with more self-reflection iterations, the performance of HybGRAG improves further. Performance improves significantly when increasing the number of iterations from $1$ to $2$, where no self-reflection is performed in iteration $1$. It is also shown that a few iterations are sufficient, as the improvement diminishes over iterations.
Model Size
Although we do not have access to Claude 3 Opus, we conduct experiments with Claude 3 Haiku, a more cost-efficient but less powerful alternative to Claude 3 Sonnet [^2] https://www.anthropic.com/news/claude-3-family. In Table 4, HybGRAG maintains strong performance even with Claude 3 Haiku. The results also follow the scaling law of LLMs (Kaplan et al., 2020).
@@ 4.2 End-to-End RAG Evaluation on CRAG
In Table 7, HybGRAG outperforms all baselines in CRAG. RAGs with a single retrieval module cannot handle both types of questions. RAG with a concatenated reference also distracts by irrelevant content in the long reference. Although the same retriever bank is provided, self-reflective baselines still find it difficult to refine their action. Since ReAct relies on the LLM’s ability to think and provide natural language feedback, it often lacks clear guidance for improving its actions. Without a fine-tuned retrieval evaluator, Corrective RAG cannot effectively identify the usefulness of a reference. This demonstrates the advantages of our critic module with corrective feedback. Furthermore, HybGRAG is robust to the choice of LLM base models. The ablation study is provided in Appx. B.2.
@@ 4.3 Model Cost Analysis
We report the number of API calls and token consumption (excluding references) for each step in an iteration of HybGRAG in Table 4.3 and 4.3 for STaRK and CRAG, respectively. While most of the token consumption arises from the examples used for ICL, the prompts themselves require very few tokens. Moreover, since HybGRAG uses the chat LLM as the router, the examples for ICL only need to be given once. Compared to the state-of-the-art baseline AvaTaR in STaRK, which requires at least $500$ API calls during training, our hybrid retrieval module achieves a relative improvement $24$% in Hit@1 with only $2$ API calls, while HybGRAG achieves $51$% with at most $14$ API calls, both without training.
@@
Commentor 1 52 1215 1583
Table 9: Number of API calls and tokens for CRAG.
HybGRAG API Call # Token # for Token # for
Component / Iteration Prompts Examples
Router 4 266 5752
Validator 1 56 0
Commentor 1 78 598
Generator 2 168 553
@@ 5 Related Works
Graph RAG (GRAG)
Various settings have been explored for GRAG (Peng et al., 2024), and can be roughly divided into three directions. The first focuses on KBQA, taking advantage of the LLM capability (Yasunaga et al., 2021; Sun et al., 2024; Jin et al., 2024; Mavromatis and Karypis, 2024). The second focuses on ODQA, building relationships between documents to improve retrieval (Li et al., 2024a; Dong et al., 2024; Edge et al., 2024). The last assumes that a subgraph is given when answering a question (He et al., 2024; Hu et al., 2024). In contrast, this paper focuses on solving HQA in SKB, and previous GRAG methods are not easily generalized to HQA.
Self-Reflective LLMs
For complex tasks, LLMs are unlikely to generate the correct output on their first attempt. Self-reflection addresses this issue by optimizing the output through a feedback-driven reflection process. A critic is commonly used to give feedback, implemented with various approaches: pre-trained LLMs (Yao et al., 2023; Shinn et al., 2023; Madaan et al., 2023), external tools (Gou et al., 2024), or fine-tuned LLMs (Paul et al., 2024; Asai et al., 2024; Yan et al., 2024). AvaTaR (Wu et al., 2024a) is the most recent work on optimizing the prompt iteratively through contrastive reasoning. In our problem, while external tools and labels for fine-tuning are not available, using pre-trained LLMs as critics requires careful designs. For example, ReAct (Yao et al., 2023) relies on the LLM’s ability to think and provide natural language feedback, which is often too implicit to support effective self-reflection in HQA.
@@ 6 Conclusions
To solve hybrid question answering (HQA), we propose HybGRAG, which is based on insights from our empirical analysis. In summary, HybGRAG has following advantages:

### v2

@@ 1 Introduction
Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Guu et al., 2020) enables large language models (LLMs) to access the information from an unstructured document database. This allows LLMs to address unknown facts and solve Open-Domain Question Answering (ODQA) with additional textual information. Graph RAG (GRAG) has extended this concept by retrieving information from structured knowledge bases, where documents are interconnected by relationships. Existing GRAG methods focus on two directions: extracting relational information from knowledge graphs (KGs) and leveraging LLMs for Knowledge Base Question Answering (KBQA) (Yasunaga et al., 2021; Sun et al., 2024; Jin et al., 2024; Mavromatis and Karypis, 2024), and building relationships between documents in the database to improve ODQA performance (Li et al., 2024a; Dong et al., 2024; Edge et al., 2024).
@@ 3. Interpretable ?
As shown in Table 1, HybGRAG is the only work that satisfies all the properties and solves HQA in SKB. As shown in Fig. 2, when evaluated on the HQA benchmark STaRK, HybGRAG outperforms the second-best baseline, achieving relative improvements in Hit@$1$ of 47% on STaRK-MAG and 55% on STaRK-Prime, respectively.
@@ 2 Proposed Insights: Challenges in HQA
What new challenges in HQA over SKB remain unsolved by existing methods? In this section, we first define the problem HQA, and then conduct experiments to uncover two critical insights, laying the foundation for designing our method for HQA.
@@ 2.1 Problem Definition
A semi-structured knowledge base (SKB) consists of a KG $G=(\mathcal{E},\mathcal{R})$, where $\mathcal{E}$ and $\mathcal{R}$ represent the sets of entities and relations. It also includes a set of text documents $\mathcal{D}=\bigcup_{i\in\mathcal{E}}{D_{i}}$, where $D_{i}$ is the document associated with entity $i$. Entity and relation types are denoted by $\mathcal{T}_{E}$ and $\mathcal{T}_{R}$, respectively. Each hybrid question $q$ in SKB involves semi-structured information, namely, textual and relational information. We define hybrid question answering (HQA) as follows:
@@ 2.2 C1: Hybrid-Sourcing Question
To investigate if it is necessary to leverage both textual and relational information to answer hybrid questions, we conduct an experiment to show that text documents and KG contain useful but non-overlapping information. As a retriever that uses only textual information, vector similarity search (VSS) (Karpukhin et al., 2020) performs retrieval and ranking by comparing the question and documents in the embedding space (“ada-002”); as a retriever that uses only relational information, Personalized PageRank (PPR) (Andersen et al., 2006) performs random walks from the topic entities identified by an LLM (Claude 3 Sonnet) and ranks neighboring entities based on their connectivity in KG.
@@ 3.1 Retriever Bank (for C1)
The hybrid retrieval module takes the identified topic entities $\hat{\mathcal{E}}$ and useful relations $\hat{\mathcal{R}}$ as input. It uses a graph retriever to extract entities in the ego-graph of $\hat{\mathcal{E}}$, connected by $\hat{\mathcal{R}}$. For example, in Fig. 1, {^=$\{\hat{\mathcal{E}}=\{\text{John Smith}\},\hat{\mathcal{R}}=\{\text{author writes paper}\}\}$ and the graph retriever extracts the entities/papers connected by the path “John Smith -> author writes paper -> {paper(s)}”. If more than one ego-graph is extracted, their intersection is used as the final result. Finally, to solve hybrid questions, we propose ranking the documents associated with the extracted entities using a VSS ranker between question $q$ and documents $\mathcal{D}$. This ensures the synergy between the relational and textual information.
@@ 3.2 Critic Module (for C2)
To solve this, we propose the critic module, which provides feedback to help the router perform better question routing. Instead of using a single LLM to complete this complicated task, we divide the critic into two parts, an LLM validator $C_{val}$ to validate the correctness of the retrieval $\mathcal{X}$, and an LLM commenter $C_{com}$ to provide feedback $f_{t}$ if the retrieval is incorrect. This divide-and-conquer step, similar to previous works (Gao et al., 2022; Asai et al., 2024), is crucial to our critic module, offering two key advantages: (1) By breaking a difficult task into two easier ones, we can now leverage pre-trained LLMs to solve them while maintaining good performance. This resolves the issue when the labels are not available for fine-tuning an LLM critic. (2) Since the tasks of $C_{val}$ and $C_{com}$ are independent, they can each have their own exclusive contexts, preventing the inclusion of irrelevant information and avoiding the “lost in the middle” phenomenon (Shi et al., 2023; Liu et al., 2024).
@@
Commenter
The LLM commenter $C_{com}$ aims to provide feedback $f$ to help the router refine question routing. To effectively guide the router, we construct corrective feedback that it can easily understand. In more detail, it points out the error(s) in each action, such as incorrect identification of topic entities, as shown in Table 4. Unlike natural language feedback, which may cause uncertainty or inconsistency depending on the LLM used, our corrective feedback provides clear guidance on how to refine the question routing. Furthermore, it leverages in-context learning (ICL) to provide sophisticated feedback. We collect a small number of successful experiences ($\approx 30$) in the training set as examples, with each experience $\{s_{t},\hat{\mathcal{E}}_{t},\hat{\mathcal{R}}_{t},f_{t+1}\}$ comprising a pair of router action and feedback, which is verified by ground truth. During inference, the commenter gives high-quality feedback based on multiple pre-collected examples.
@@ 3.3 Overall Algorithm
The algorithm of HybGRAG is in Algo. 1. Given a question $q$, in iteration $t$, the router determines $s_{t}$, ^$\hat{\mathcal{E}}_{t}$ and ^$\hat{\mathcal{R}}_{t}$ to retrieve the references $\mathcal{X}_{t}$ from both $G$ and $\mathcal{D}$ in SKB, or only $\mathcal{D}$, with the selected retrieval module. The validator $C_{val}$ in the critic module then decides whether to accept $\mathcal{X}_{t}$ as the final answer or reject it. If $\mathcal{X}_{t}$ is rejected, the commenter $C_{com}$ generates feedback $f_{t+1}$ for the router to assist in refining its action in iteration $t+1$.
@@ 4 Experiments
    Effectiveness: How well does HybGRAG perform in real-world benchmarks?
@@
Figure 3: Design choices in HybGRAG are necessary in STaRK. We compare HybGRAG with two variants: a validator without validation context, and a commenter with only 5-shot. Oracle uses ground truth during inference.
@@
Table 7: HybGRAG performs best with multi-agent design in STaRK-MAG. “Router for SR” baseline performs self-reflection using a single LLM router.
Method Setting Hit@1 Hit@5 Recall@20 MRR
Hybrid RM No-Agent 0.5028 0.5820 0.5031 0.5373
Router for SR Single-Agent 0.6206 0.7069 0.6187 0.6587
HybGRAG Multi-Agent 0.6540 0.7531 0.6570 0.6980
Benchmarks
We conduct experiments on two QA benchmarks: STaRK  [^1] Due to legal issue, only STaRK-MAG and STaRK-Prime are used in this article. (Wu et al., 2024b), which serves as the primary evaluation benchmark and focuses on retrieval, and CRAG (Yang et al., 2024), a complementary benchmark to evaluate end-to-end RAG performance. While STaRK focuses on HQA, CRAG encompasses both ODQA and KBQA as sub-problems. Detailed benchmark descriptions are provided in Appx. A.
@@ 4.1.2 Ablation Study (RQ2)
Critic Module We compare HybGRAG variants with a validator without validation context, a commenter with only five shots, and those with oracles. The oracle has access to the ground truth, which gives the optimal judgement on the correctness of the output and the error type of question routing, if there is any. In Fig. 3, we show that HybGRAG performs the best with all our design choices, approaching the performance of an oracle.
Self-Reflection In Fig. 4, we demonstrate that with more self-reflection iterations, the performance of HybGRAG improves further. Performance improves significantly when increasing the number of iterations from $1$ to $2$, where no self-reflection is performed in iteration $1$. It is also shown that a few iterations are sufficient, as the improvement diminishes over iterations.
Model Size Although we do not have access to Claude 3 Opus, we conduct experiments with Claude 3 Haiku, a more cost-efficient but less powerful alternative to Claude 3 Sonnet [^2] https://www.anthropic.com/news/claude-3-family. In Table 4, HybGRAG maintains strong performance even with Claude 3 Haiku. The results also follow the scaling law of LLMs (Kaplan et al., 2020).
Multi-Agent Perspective Since HybGRAG can be interpreted as a multi-agent system, we add a single-agent baseline, which relies on the router to make decisions and provide feedback for self-reflection. In Table 4, HybGRAG outperforms both single-agent and no-agent baselines. This highlights that self-reflection is essential for achieving strong performance in HQA, as pointed out in Challenge 2. Moreover, unlike the plain text feedback generated by the single-agent baseline, the feedback generated by HybGRAG more effectively guides the router in refining its decision, thanks to our carefully designed critic module.
@@ 4.2 End-to-End RAG Evaluation on CRAG
In Table 8, HybGRAG outperforms all baselines in CRAG. RAGs with a single retrieval module cannot handle both types of questions. RAG with a concatenated reference also distracts by irrelevant content in the long reference. Although our retriever bank is provided, agentic and self-reflective baselines still struggle to refine their actions. Since ReAct relies on the LLM’s ability to think and provide natural language feedback, it often lacks clear guidance. Without a fine-tuned retrieval evaluator, Corrective RAG cannot effectively identify the usefulness of a reference.
@@ 4.3 Model Cost Analysis
We report the number of API calls and token consumption (excluding references) for each step in an iteration of HybGRAG in Table 4.3 and C.2 for STaRK and CRAG, respectively. While most of the token consumption arises from the examples used for ICL, the prompts themselves require very few tokens. Moreover, since HybGRAG uses the chat LLM as the router, the examples for ICL only need to be given once. Compared to the state-of-the-art baseline AvaTaR in STaRK, which requires at least $500$ API calls during training, our hybrid retrieval module achieves a relative improvement $24$% in Hit@1 with only $2$ API calls, while HybGRAG achieves $51$% with at most $14$ API calls, both without training.
@@
Commenter 1 52 1215 1583
@@ 5 Related Works
Graph RAG (GRAG) Various settings have been explored for GRAG (Peng et al., 2024), and can be roughly divided into three directions. The first focuses on KBQA, taking advantage of the LLM capability (Yasunaga et al., 2021; Sun et al., 2024; Jin et al., 2024; Mavromatis and Karypis, 2024). The second focuses on ODQA, building relationships between documents to improve retrieval (Li et al., 2024a; Dong et al., 2024; Edge et al., 2024). The last assumes that a subgraph is given when answering a question (He et al., 2024; Hu et al., 2024). In contrast, this paper focuses on solving HQA in SKB, and previous GRAG methods are not easily generalized to HQA.
Agentic and Self-Reflective LLMs LLM agents (Yao et al., 2023; Wu et al., 2024a) facilitate planning in complex reasoning tasks. Among them, AvaTaR is the most recent, proposing iterative prompt optimization via contrastive reasoning. However, they may still struggle to generate the correct output on the first attempt. Self-reflection addresses this limitation by iteratively optimizing the output based on feedback, typically provided by a critic implemented using various approaches: pre-trained LLMs (Shinn et al., 2023; Madaan et al., 2023), external tools (Gou et al., 2024; Qiao et al., 2024), or fine-tuned LLMs (Paul et al., 2024; Asai et al., 2024; Yan et al., 2024). Nevertheless, they do not generalize to HQA for two reasons. First, they lack appropriate retrieval tools and guidance on how to refine retrieval effectively. Second, in the absence of external tools or labels for fine-tuning, using pre-trained LLMs as critics without careful design results in suboptimal self-evaluation and overly implicit feedback.
@@ 6 Conclusions
To solve hybrid question answering (HQA), we propose HybGRAG, driven by insights from our empirical analysis, which has following advantages:
@@
Limitations
While HybGRAG is capable of outperforming existing RAG and GRAG methods on HQA, it still has some limitations: (1) HybGRAG uses only the simplest retrieval modules, and various alternatives are not explored. For example, the ranker in the retrieval modules could be replaced with a cross-encoder ranker, and the retriever in the hybrid retrieval module could use the top-$K$ entities from PPR instead. (2) HybGRAG does not offer significant advantages in terms of domain adaptation. In experiments, although HybGRAG outperforms baselines, its performance on STaRK-Prime is worse than in STaRK-MAG, where the academic domain is generally considered less complex than the medicinal domain. (3) The commentor in HybGRAG selects random experiences when performing ICL. For example, selecting experiences with questions most relevant to the current one may yield better performance. Although these limitations point out areas for potential improvement, they also present future directions to further enhance the capabilities of HybGRAG.
@@ C.2 Prompts
Table 14: Number of API calls and tokens for CRAG.
HybGRAG API Call # Token # for Token # for
Component / Iteration Prompts Examples
Router 4 266 5752
Validator 1 56 0
Commenter 1 78 598
Generator 2 168 553